# LBM: Latent Bridge Matching for Fast Image-to-Image Translation

### v1

@@ Abstract
In this paper, we introduce Latent Bridge Matching (LBM), a new, versatile and scalable method that relies on Bridge Matching in a latent space to achieve fast image-to-image translation. We show that the method can reach state-of-the-art results for various image-to-image tasks using only a single inference step. In addition to its efficiency, we also demonstrate the versatility of the method across different image translation tasks such as object removal, normal and depth estimation, and object relighting. We also derive a conditional framework of LBM and demonstrate its effectiveness by tackling the tasks of controllable image relighting and shadow generation. We provide an open-source implementation of the method at https://github.com/gojasper/LBM.
@@ 4.1 Object-removal
m.$\downarrow$ $11.25^{\circ}\uparrow$ $30.0^{\circ}\uparrow$ m.$\downarrow$ $11.25^{\circ}\uparrow$ $30.0^{\circ}\uparrow$ m.$\downarrow$ $11.25^{\circ}\uparrow$ $30.0^{\circ}\uparrow$ m.$\downarrow$ $11.25^{\circ}\uparrow$ $30.0^{\circ}\uparrow$ Rank
@@ A.1 Object-removal task
For the object-removal task, we trained our model for 20k iterations on 2 H100 GPUs. We set $\sigma=0.05$ and used the timestep distribution we propose in the main paper i.e. $\pi(t)=\mathcal{U}({i/4})_{i\in\{{0,1,2,3}\}}$. We use a bucketing strategy as proposed in [66] allowing us to handle multiple aspect ratios and resolutions. This strategy consists of defining buckets with pre-defined aspect ratios and pixel budgets and filling them with the data flow. During each training iteration, a target pixel budget is sampled and then the upcoming images are assigned to the bucket with the closest aspect ratio and budget and are resized accordingly. We use the following bucket pixel budgets: $[256^{2},512^{2},768^{2},1024^{2}]$ sampled with probabilities $[0.1,0.2,0.2,0.5]$. For each budget we consider aspect ratios ranging from $0.25$ to $4$. The batch sizes are respectively set to 32, 16, 8 and 4 for each budget. We trained the model with LPIPS pixel loss with weight $\lambda=10$ and a learning rate of $3e^{-5}$ and we used the AdamW optimizer [52]. For data sources, we randomly sampled data from the RORD train set, our synthetic dataset or our in-the-wild dataset with probabilities $[0.3,0.3,0.4]$. For the latter, we used the random masking strategy proposed in [89] while for RORD and our synthetic dataset we used the provided semantic masks. The denoiser is initialized using the weights of the pre-trained text-to-image model SDXL [66].
@@ Appendix B Additional object-removal results
Table 5: Metrics for object-removal task computed on RORD validation set (52k images) using the fine semantic masks. Our method uses a single NFE. Best results are highlighted in bold, second best are underlined.
@@ Appendix C Results for depth estimation
DiverseDepth 11.7 87.5 - 19.0 70.4 - 22.8 69.4 - 10.9 88.2 - 37.6 63.1 - 15.6
MiDaS 11.1 88.5 - 23.6 63.0 - 18.4 75.2 - 12.1 84.6 - 33.2 71.5 - 15.1
LeRes 9.0 91.6 - 14.9 78.4 - 17.1 77.7 - 9.1 91.7 - 27.1 76.6 - 12.2
Omnidata 7.4 94.5 - 14.9 83.5 - 16.6 77.8 - 7.5 93.6 - 33.9 74.2 - 12.2
DPT 9.8 90.3 - 10.0 90.1 - 7.8 94.6 - 8.2 93.4 - 18.2 75.8 - 9.9
HDN 6.9 94.8 - 11.5 86.7 - 12.1 83.3 - 8.0 93.9 - 24.6 78.0 - 9.2
GenPercept 5.6 96.0 99.2 13.0 84.2 97.2 7.0 95.6 98.8 6.2 96.1 99.1 35.7 75.6 86.6 7.9
Diff.-E2E-FT 5.4 96.5 99.1 9.6 92.1 98.0 6.4 95.9 98.7 5.8 96.5 98.8 30.3 77.6 87.9 5.4
DepthAnything V2 4.5 97.9 99.3 7.4 94.6 98.6 13.1 86.5 97.5 4.2 97.8 99.3 26.5 73.4 87.1 5.1
DepthAnything 4.3 98.1 99.6 7.6 94.7 99.2 12.7 88.2 98.3 4.3 98.1 99.6 26.0 75.9 87.5 3.9
GeoWizard 5.6 96.3 99.1 14.4 82.0 96.6 6.6 95.8 98.4 6.4 95.0 98.4 33.5 72.3 86.5 9.3
Marigold (LCM) 6.1 95.8 99.0 9.8 91.8 98.7 6.8 95.6 99.0 6.9 94.6 98.6 30.7 77.5 89.3 7.2
Marigold 5.5 96.4 99.1 9.9 91.6 98.7 6.5 95.9 99.0 6.4 95.2 98.8 30.8 77.3 88.7 6.2
Lotus-D 5.1 97.2 99.2 8.1 93.1 98.7 6.1 97.0 99.1 5.5 96.5 99.0 22.8 73.8 86.2 3.9
Lotus-G 5.4 96.8 99.2 8.5 92.2 98.4 5.9 97.0 99.2 5.9 95.7 98.8 22.9 72.9 86.0 5.1
Ours 5.6 97.2 99.2 9.4 93.0 98.9 6.3 96.5 99.3 5.7 97.0 99.2 30.3 77.5 89.3 3.5
@@ D.2 Image relighting
Appendix E Additional samples

### v2

@@ Abstract
In this paper, we introduce Latent Bridge Matching (LBM), a new, versatile and scalable method that relies on Bridge Matching in a latent space to achieve fast image-to-image translation. We show that the method can reach state-of-the-art results for various image-to-image tasks using only a single inference step. In addition to its efficiency, we also demonstrate the versatility of the method across different image translation tasks such as object removal, normal and depth estimation, and object relighting. We also derive a conditional framework of LBM and demonstrate its effectiveness by tackling the tasks of controllable image relighting and shadow generation. We provide an implementation at https://github.com/gojasper/LBM.
@@ 4.1 Object-removal
[…]
@@ A.1 Object-removal task
For the object-removal task, we trained our model for 20k iterations on 2 H100 GPUs. We set $\sigma=0.05$ and used the timestep distribution we propose in the main paper i.e. $\pi(t)=\mathcal{U}({i/4})_{i\in\{{0,1,2,3}\}}$. We use a bucketing strategy as proposed in [66] allowing us to handle multiple aspect ratios and resolutions. This strategy consists of defining buckets with pre-defined aspect ratios and pixel budgets and filling them with the data flow. During each training iteration, a target pixel budget is sampled and then the upcoming images are assigned to the bucket with the closest aspect ratio and budget and are resized accordingly. We use the following bucket pixel budgets: $256^{2},512^{2},768^{2},1024^{2}]$ sampled with probabilities $[0.1,0.2,0.2,0.5]$. For each budget we consider aspect ratios ranging from $0.25$ to $4$. The batch sizes are respectively set to 32, 16, 8 and 4 for each budget. We trained the model with LPIPS pixel loss with weight $\lambda=10$ and a learning rate of $3e^{-5}$ and we used the AdamW optimizer [52]. For data sources, we randomly sampled data from the RORD train set, our synthetic dataset or our in-the-wild dataset with probabilities $[0.3,0.3,0.4]$. For the latter, we used the random masking strategy proposed in [89] while for RORD and our synthetic dataset we used the provided semantic masks. The denoiser is initialized using the weights of the pre-trained text-to-image model SDXL [66].
@@ Appendix B Additional object-removal results
Table 5: Metrics for object-removal task with models fine-tuned on RORD train set and evaluated on RORD validation set (52k images) using the fine semantic masks. Our method uses a single NFE. Best results are highlighted in bold, second best are underlined.
For the sake of completeness, we also fine-tune LAMA, SDXL-inpaint., PowerPaint and our LBM checkpoint (Attentive Eraser is training-free) only on RORD train set such that all the models see approx. 400k samples, which was enough to reach convergence. For the sake of completeness, we also train a LBM model from scratch only on the RORD train set with the same number of iterations. We share the results in Tab. 6. As shown in the table, while this fine-tuning step improves competitors’ results, in particular for fine masks, our method still outperforms competitors for most metrics. Also note that our initial model is 047 trained on 2 H100 for $\approx$18h vs. 240h on 8 V100 for LAMA.
[…]
Coa. Fin. Coa. Fin. Coa. Fin. Coa. Fin. Coa. Fin. time (s)
LAMA 30.3 21.4 38.0 28.2 1592.2 1350.3 19.7 20.6 55.9 57.1 0.1
SDXL-inp. 27.2 18.5 27.3 18.0 2297.3 2213.1 19.8 21.4 64.9 69.0 7.2
PowerPaint 29.9 27.0 30.0 23.7 2871.2 2679.7 18.5 19.9 58.3 63.4 4.2
AE 29.7 18.4 33.2 22.2 2029.0 1773.0 20.9 22.8 65.7 70.8 8.0
Ours 26.9 15.7 30.5 15.6 1306.6 997.4 22.5 24.5 69.2 73.2 0.3
Ours (scratch) 27.9 16.7 30.7 16.9 1329.5 1032.2 22.4 24.4 69.0 72.9 0.3
Table 6: Metrics for object-removal task computed on RORD validation set using the coarse (Coa.) and fine (Fin.) masks. Our method and LAMA use a single neural function evaluation (NFE), others use 50 NFEs. Inference time is averaged over 50 images and computed on a single H100 GPU.
@@ Appendix C Results for depth estimation
DiverseDepth 11.7 87.5 - 19.0 70.4 - 22.8 69.4 - 10.9 88.2 - 37.6 63.1 - 16.6
MiDaS 11.1 88.5 - 23.6 63.0 - 18.4 75.2 - 12.1 84.6 - 33.2 71.5 - 16.1
LeRes 9.0 91.6 - 14.9 78.4 - 17.1 77.7 - 9.1 91.7 - 27.1 76.6 - 13.2
Omnidata 7.4 94.5 - 14.9 83.5 - 16.6 77.8 - 7.5 93.6 - 33.9 74.2 - 13.2
DPT 9.8 90.3 - 10.0 90.1 - 7.8 94.6 - 8.2 93.4 - 18.2 75.8 - 10.8
HDN 6.9 94.8 - 11.5 86.7 - 12.1 83.3 - 8.0 93.9 - 24.6 78.0 - 10.2
DepthFM 6.0 95.5 - 9.1 90.2 - 6.5 95.4 - 6.6 94.9 - 22.4 78.5 - 7.2
GenPercept 5.6 96.0 99.2 13.0 84.2 97.2 7.0 95.6 98.8 6.2 96.1 99.1 35.7 75.6 86.6 8.3
Diff.-E2E-FT 5.4 96.5 99.1 9.6 92.1 98.0 6.4 95.9 98.7 5.8 96.5 98.8 30.3 77.6 87.9 5.6
DepthAnything V2 4.5 97.9 99.3 7.4 94.6 98.6 13.1 86.5 97.5 4.2 97.8 99.3 26.5 73.4 87.1 5.4
DepthAnything 4.3 98.1 99.6 7.6 94.7 99.2 12.7 88.2 98.3 4.3 98.1 99.6 26.0 75.9 87.5 4.1
GeoWizard 5.6 96.3 99.1 14.4 82.0 96.6 6.6 95.8 98.4 6.4 95.0 98.4 33.5 72.3 86.5 9.6
Marigold (LCM) 6.1 95.8 99.0 9.8 91.8 98.7 6.8 95.6 99.0 6.9 94.6 98.6 30.7 77.5 89.3 7.7
Marigold 5.5 96.4 99.1 9.9 91.6 98.7 6.5 95.9 99.0 6.4 95.2 98.8 30.8 77.3 88.7 6.4
Lotus-D 5.1 97.2 99.2 8.1 93.1 98.7 6.1 97.0 99.1 5.5 96.5 99.0 22.8 73.8 86.2 4.0
Lotus-G 5.4 96.8 99.2 8.5 92.2 98.4 5.9 97.0 99.2 5.9 95.7 98.8 22.9 72.9 86.0 5.3
Ours 5.6 97.2 99.2 9.4 93.0 98.9 6.3 96.5 99.3 5.7 97.0 99.2 30.3 77.5 89.3 3.7
@@ D.2 Image relighting
Appendix E Memory footprint and inference time
Our intuition to use a latent model is motivated by the key observations made in [73] where the authors scale image generation from diffusion models. Nevertheless, we quantitatively report in Tab. 8 the memory/latency comparison between a pixel model and a latent model both for training and inference. Note that the VAE compresses the source image by a factor of 8 and is frozen during training drastically reducing the memory footprint of the model as shown in the table.
Mode (Resolution) Metric Pixel Model Latent Model Gain
Inference (256 / 1024) Latency (s) 0.19 / 20.11 0.14 / 0.27 26.3% / 98.7%
Peak Memory (Gb) 5.63 / 15.71 5.29 / 7.70 6.0% / 51.0%
Training (256 / 1024) Latency (s) 0.71 / - 0.43 / 0.58 39.4% / -
Peak Memory (Gb) 43.19 / OoM 24.85/ 25.35 41.3% / -
Table 8: Training and inference memory usage and per-iteration latency for a pixel and a latent bridge model. The metrics are averaged over 10 images using a batch size of 1 with AdamW for training and 1 NFE for inference on a single H100 80Gb GPU.
Appendix F Additional samples